Day shift 6/9/99 summary

Valeri Jejer (jejer@fnal.gov)
Wed, 9 Jun 1999 15:35:56 -0500 (CDT)

Shift Summary for day shift, 09/06/99

On Shift: Valeri J., Greg B.

Also : Ed, Sydney, Elizabeth, Sandrine, Val, Miguel, Jasons, Jianbo,
Ricks F. & C.

Hours of beam delivered at all intensities:
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 13463 - 3.05M events - BEAM_832
Run 13464 - 2.6M events - BEAM_832
Run 13465 - 0 events - COSMICS

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
7:20 - noon target scan
noon - beam interlock studies

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13463 38-72 7:00 Noted that MU counter illumination plots don't look
right because weeping magnets had wrong settings:
NM2S2 was set to 500 A instead of 1500 A and NM2S3
to 100 A instead of 500 A. Sasha L. (owl shift)
called MCR and asked to set the magnets to nominal
values. Rick F. found that those settings were on
from 6:20 to 7:05, and why is unknown. SDAQ log
shows the magnet currents out of limits as well.
DC currents were OK.

82 7:20 HCC busy -> kfrend.

7:53 Stop run 13463.

13464 7:59 Start run 13464 (BEAM_832), target scan was going
on through this run.

14 8:20 pipeline errors -> reset pipeline w/DBC -> kfrend

8:45 pipeline errors -> reset pipeline w/DBC -> kfrend

8:52 lost beam

9:40 beam back

114 10:32 pipeline errors -> reset pipeline w/DBC -> kfrend

11:41 lost beam

12:10 RC, RF and GB finished target scans. Beam is
centered on target. 90 degrees monitor has large
fluctuations, maybe because of spill quality or
alignment of 1/8"x1/8" hole. RF went to MCR to do
beam interlock and quad studies.

12:11 Ch.13: "KTeV beam interlock studies in progress".
Beam is of very low intensity and unknown quality.

12:43 Stop run 13464

14:10 Still no beam, start cosmic run 13465, Jason H.
investigates HCC FULL problem.

14:35 Stop run 13465.

14:40 RF begins quad studies, estimate - 1-2 hours.

14:58 Ch.13: "We are making access into Booster and SY to
check for a large LCW leak. No est. at this time".

15:05 Start cosmic run 13465 since there is nothing better
to run.
========================================================================
Follow up Notes on Previous Shift

CSI 1847 has been fine (EB).

========================================================================
Items to Follow up on Next Shift